(NOS. 2-5) PROBLEM 1. A cylindrical container of height equal to twice the diameter of its base can hold 12 liters (1L= 1,000 cm3) of water. Another cylindrical container with the same capacity has its height equal to three times the diameter of its base.



_______________________ 4) Determine the amount of aluminum required for making the first
container in cm2?


_____________________ 5) Determine the amount of aluminum required for making the
second container in cm2?
1
Expert's answer
2020-11-30T20:52:53-0500

 The amount of aluminum required for making the first container in cm2

Solution

We, first of all calculate the diameter of the first container.

We know that the volume of a cylinder is given by, V=πr2h, where r = radius of the base surface.

Diameter = 2 × radius

Therefore, radius (r) = d​/2

In this case, height = twice the diameter of the base

Therefore, Volume = π (d/2)2​×h= 12 × 1000cm3

π (d2​)/4 × 2d=12×1000cm3, where π =22/7

We then solve for d

d = [(12000×2)/π]​1/3 = 19.69cm​

Diameter = 19.69cm

To calculate, the amount of aluminum required, we assume that the containers are closed.

Amount of aluminum = total surface area = 2πrh + 2 πr2

Where radius = 19.69/2 = 9.85cm

Height = 19.69 × 2 = 39.38 cm.

Therefore, amount of aluminum for the first container = (2 × π × 9.85cm × 39.38cm) + [(2 × π × (9.85cm) 2] = 3048.04cm3

The amount of aluminum required for making the second container in cm2

Solution

Let d1​= the diameter of the base of the first container, h1​= the height of the first container and V1​= the volume of the first container.

Similarly, let d2​= the diameter of the base of the second container, h2​= the height of the second container and V2​= the volume of the second container.

Therefore, from the given information, h1​=2d1, V1​=π (d1​​/2)2, h1​=1/2​π d13=1/16πh31

h2=3d2, V1 = π (d2​​/2)2, h2​= 3/4π d23=1/36πh32

We are given that, V1​=V2​=12L=12000cm3

​Therefore, 3/4π d23=12,000cm3  

Calculate the value of d2

d2 = [(4 × 12000)/3π] 1/3

d2 = 17.20cm

The diameter of second container = 17.20cm.

Therefore, radius of second container = 17.20/2 = 8.60cm

Height of second container = three times the diameter of the base = 3 × 17.20 = 51.60cm

Hence, the amount of aluminum required to make second container = total surface area = (2 × π × 8.60 × 51.60) + [2 × π × (8.6)2] = 3254.24cm2